1) A random sample is drawn from a population of known standard deviation 10. Construct a 90% confidence interval for the population mean based on the information given n = 36 š‘„ = 100


1
Expert's answer
2022-05-17T17:40:00-0400

The critical value forĀ "\alpha = 0.1"Ā isĀ "z_c = z_{1-\alpha/2} = 1.6449."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-z_c\dfrac{\sigma}{\sqrt{n}}, \bar{X}+z_c\dfrac{\sigma}{\sqrt{n}})"

"=(100-1.6449\dfrac{10}{\sqrt{36}}, \bar{X}+z_c\dfrac{\sigma}{\sqrt{n}})"

"=(97.2585, 102.7415)"

Therefore, based on the data provided, theĀ 90%Ā confidence interval for the population mean isĀ "97.2585 < \mu < 102.7415," which indicates that we areĀ 90%Ā confident that the true population meanĀ "\mu" is contained by the intervalĀ "(97.2585, 102.7415)."